OpenStudy (anonymous):

A deposit of Po into a bank account has a doubling time of 50 years. No other deposits or withdrawals are made. a. How much money is in the bank account after 50 years? 100 years? 150 years? b. How many times does the amount of money double in t years? Use this to write a formula for P1, the amount of money in the account after t years. ***I'm confused on how to do this... Could you please explain this to me? Thank you :)

OpenStudy (loser66):

a) \(P_0, 2P_0,4P_0,8P_0\)

OpenStudy (anonymous):

sorry i'm not sure what that means? :/

OpenStudy (loser66):

from yours: A deposit of Po into a bank account has a doubling time of 50 years At the beginning, the amount of money is Po right? after 50 years, it is double, it means at that time, the amount of money is 2Po

OpenStudy (anonymous):

ohh i see.. wait was that supposed to be latex? this is what shows up on my side :/

OpenStudy (loser66):

I don't know why, I cannot use draw box below, either.

OpenStudy (anonymous):

so for a, the answers should be: for 50 years: 2Po for 100 years: 4Po for 150 years: 8Po ? yeah OS has been weird lately haha :/

OpenStudy (loser66):

yes, that's what I mean

OpenStudy (anonymous):

okay awesome!! :) so are we on part b now? :)

OpenStudy (loser66):

There is someone watch us for free; I think now is the time for his turn, @eliassaab

OpenStudy (anonymous):

haha okay :P

OpenStudy (anonymous):

Continue, you are doing fine so far

OpenStudy (anonymous):

how do we start part b? :/

OpenStudy (anonymous):

\[\Large P(t) = P_0 e^{\frac{t \ln 2}{50}} \]

OpenStudy (anonymous):

not sure why, but latex isn't showing on my end right now :/ I can't tell what that says... would you be able to type it out? :/ that's what it looks like on my end :/ OS is acting up again haha

OpenStudy (anonymous):

@eliassaab @Loser66 just wondering, can you see what I drew here?

OpenStudy (anonymous):

oh wow... i drew and it didn't even show up lol :/

OpenStudy (anonymous):

Hopefully that'll be fixed soon!!

OpenStudy (anonymous):

Insert this P(t) = P_0 e^{\frac{t \ln 2}{50}} in http://www.codecogs.com/eqneditor

OpenStudy (anonymous):

wait, so which equation is it? @eliassaab i got this when I put it in :/

OpenStudy (anonymous):

and @Loser66 this is what showed up on your doc.. haha which one is it? :/

OpenStudy (anonymous):

That is the formula

OpenStudy (loser66):

hahahah, hand off, OS said that: " hey!! you guys work very hard, now take a rest. friends" I obey its instruction

OpenStudy (anonymous):

lol :p okay... wait @eliassaab which one is it? yours? or @Loser66 ??

OpenStudy (loser66):

me done, follow @eliassaab

OpenStudy (anonymous):

P(t) = P0 e^( (t ln(2)/50)

OpenStudy (anonymous):

haha okay thanks for helping @Loser66 :)

OpenStudy (anonymous):

so it's this one?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay awesome!! thank you sooo much!! :) hopefully latex and draw will work again soon on OS haha :P thanks again!! :)